The free-fall acceleration on the surface of Jupiter, the most massive planet, is 24.79 m/s^2. Jupiter's radius is 7.0
Answer #1

g= G*M/(R2)

=>24.79 = (6.67*10-11)*M/(7*107)2

=>M=1.821154423*1027   kg